OpenStudy (anonymous):

a triangular region of a community college parking lot was measured. The measure of the first angle of this triangle is double the measure of the second angle. The measure of the third angle is 131 degrees greater than double the measure of the first angle. what are the measurements of all three angles? please help!

OpenStudy (anonymous):

triangle has 180 degree: st+nd+rd=180 (2(x)) +( x) +(131)=180 2x+x+131=180 3x = 180-131 x = 49/3 x = 16.3 32.7 , 16.3, 131